Status Report

NASA Space Station Status 8 Jun 2004

By SpaceRef Editor
June 8, 2004
Filed under , , ,
NASA Space Station Status 8 Jun 2004

All ISS systems continue to function nominally except those noted previously or below.  The crew was commended on their hard work with EVA tool preparation yesterday. 

Preparations continued for the Orlan-M training on 6/11 and the subsequent EVA.   [The CDR will be wearing Orlan #25 (red markings) with BRTA radio telemetry unit #13 installed, while the FE’s suit will be #26 (blue markings) with BRTA #18.]

CDR Gennady Padalka, assisted by Mike Fincke as Crew Medical Officer (CMO), completed his second session of the MO-5 MedOps protocol of cardiovascular evaluation during graded exercises on the VELO cycle ergometer.  Mike’s turn with MO-5 is scheduled for tomorrow.   [The assessment uses the Gamma-1 ECG equipment with biomed harness, skin electrodes and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer’s instrumentation panels.  For the graded exercise, Gennady worked the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each.]

FE/SO Fincke and Padalka continued readying and preparing the U.S. EVA tools for later transfer to the Russian segment (RS) and the DC1 docking module, based on a detailed usage and stowage plan uplinked by MCC-H.   [The equipment preparations are being videotaped with the U.S. digital DVCAM for downlink via Ku-band and supported by tagup with ground specialists via S-band.  In addition to the required tool complement, the packing includes some “spare equipment” in the event that it may be needed.]         
The Orlan suits were activated, and each crewmember gathered, laid out, and installed the replaceable components (OTA) and auxiliary gear for their particular Orlan “skaphandr” suits, complete with fit checks.  The activities were supported by tagup with ground specialists.   [Assembled were portable O2 tanks (BK-3), storage batteries, LiOH canisters, moisture collectors, KVO liquid cooling garments, ShL-10 headsets, GP-10K gloves, BK-10 undergarments, socks, filters for feedwater lines (FOR), IK Orlan measurement unit and BOS degassing pump, etc.]

Orlan preparations also included installation of U.S. tools on the suits and of retaining clips on their water hoses to prevent crimping and ensure unimpeded coolant flow.  The KPU container was configured with items required for the EVA, and the BKDO contamination monitoring unit was set up to be attached near the DC-1 EVA hatch.
Later, the crew downlinked the video taken of the preparations with the U.S. DVCAM, via Ku-band.

Fincke performed the routine maintenance of the SM’s SOZh life support system.

The crew completed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED expander and VELO ergometer with load trainer. 

The ground has come up with two options for Mike Fincke’s “Saturday Science” program this weekend, for his choice if he is interested.  The options are test sequences for the Capillary Flow Experiment (CFE), which investigates fluid behavior in capillary systems, or a teleconference with the Principal Investigator (PI) for the Binary Colloidal Alloy Test-3 (BCAT-3) experiment.

Today’s CEO (Crew Earth Observations) photo targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Aral Sea (two passes today over this target area.  As ISS approached from the SW, the crew was to try the shorter settings of the variable lens for context views of the entire (shrinking) sea.  Focus should not have been an issue around 100mm.  For the second pass, light was still adequate. As the station tracked from the NW this time, the crew was again to try the shorter settings of the variable lens for those context views), Athens, Greece (pass may have been just early enough in the day to beat the clouds.  Nadir pass over the Greek capital, if attitude permitted), Istanbul, Turkey (just 82 seconds after Athens, ISS had almost the same conditions for this metropolis straddling the Bosporus channel), Red River Basin (an excellent, late morning pass in fair weather over this large target area.  Of particular interest are water levels of the Red River and especially the large man-made Lake Texoma.  Shooting for details of these features), Chicago, Illinois (a near-nadir pass of “the windy city), and The Great Salt Lake (as the decade-long drought continues in the western US, this landmark feature is experiencing dramatic shoreline and water color changes.  Attitude permitting, the crew was to try mapping details of the shoreline, especially the populous eastern side.  Most of the lake should have been just left of track).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. & Russian Segment Status (as of this morning, 1:15am EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.2; ppO2 (mmHg) — 154.2; ppCO2 (mmHg) — 3.3;
  • SM Transfer Compartment:  Pressure (mmHg) — 752; temperature (deg C) — 20.8.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 749.61; temperature (deg C) — 23.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 751.38; temperature (deg C) — 24.3; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 751.58; temperature (deg C) — 24.9; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg).
  • SM batteries:  Battery #2 is in “Cycle” mode.  All other batteries (7) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; battery #2 is in “Cycle” mode; all other batteries (4) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3372 kg (7434 lb) as of 6/3/04;  [SM(552) + FGB(2820) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:10am EDT [= epoch]):

  • Mean altitude — 362.2 km
  • Apogee — 365.9 km
  • Perigee — 358.6 km
  • Period — 91.79 min.
  • Inclination (to Equator) —  51.6329 deg
  • Eccentricity — 0.0005395
  • Solar Beta Angle —  -24.5 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 95 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 31694

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.